Net cash from operating activities $46,050 $38,000 $32,000 $16,050 shares outstanding 33,000 30,000 Compute these values and ratios for 2016 and 2017. (Round Earnings per share to 2 decimal places, a.g. $2.78 and Current Ratie and Debt to assets ratio to 1 decimal place, e.g. 1.515
31 2016 Cost of goods sold $8200810 Net income Additional information: Net cash from operating activities 46,050 $38,000 paid on common shares to 1 plecs, 15.2. compete these eaues se atostor 2016 and 2017. (Raand Earnings per share to a decinat pacte ea. s2.78 and Current Rati, and Debt to a (b) Working capital.
Assignment Gradebook ORION Downloadable eTextbook tice ment Problem 2-6A Your answer is partially correct. Try again. Condensed balance sheet and income statement data for Sheridan Company are presented as follows. Sheridan Company Balance Sheets December 31 2016 Assets Cash Recelvables (net) Other current assets Long-term investments Property, plant, and equipment (net) 2017 $29,050 21,050 76,500 87,500 60,000 524,500 484,500 804,550 729,550 84,500 104,500 62,000 Total assets Liabilities and Stockholders Equity Current liabilities Long-term liabilities Common stock s 76,050 71,050 94,500104,500 344,500 314,500 289.500239,500 $804,550 $729,550 Retained earnings Total liabilities and stockholders equity Privao bhSY ถ. 2009:2019,lahnwiex J.Sens.1nL Ali Rights Reserved. A Division of 1
Answer #1

Answer :

2016 2017
(a). Earnings per share $ 2.70 $ 2.49
(b). Working Capital $ 114,000 $ 142,000
(c). Current Ratio 2.6 : 1 2.9 : 1
(d). Debt to Assets ratio 24.1 % 21.2 %
(e). Free cash flow $ 3,000 $ 19,900

Working Notes :

(a). Earnings per share =

(Net Income - Preference shares dividend ) ÷ Weighted average shares outstanding

2016 = ( 81,000 - 0) ÷ 30,000 = $ 2.70

2017 = ( 82,000 - 0) ÷ 33,000 = $ 2.49

(b). Working Capital =

Current Assets - Current Liabilities

(Cash + Receivables+ Other current assets) - Current Liabilities

2016 = (21,050 + 76,500 + 87,500) - 71,050 = $ 114,000

2017 = (29,050 + 84,500 +104,500)- 76,050 = $ 142,000

(c). Current Ratio =

Current Assets ÷ Current Liabilities

2016 = 185,050 ÷ 71,050 = 2.6 : 1

2017 = 218,050 ÷ 76,050 = 2.9 : 1

(d). Debt to Assets ratio =

Total Liabilities ÷ Total Assets

2016 = [(71,050 + 104,500) ÷ 729,550] × 100 = 24.1 %

2017 = [(76,050 + 94,500) ÷ 804,550] × 100 = 21.2 %

(e). Free Cash Flow =

Cash flow from operating activities - Capital expenditure - Cash dividend paid

2016 = 57,050 - 38,000 - 16,050 = $ 3,000

2017 = 97,950 - 46,050 - 32,000 = $ 19,900